# CHOSEN: Contrastive Hypothesis Selection for Multi-View Depth Refinement

### v1

@@ Abstract
We propose CHOSEN, a simple yet flexible, robust and effective multi-view depth refinement framework. It can be employed in any existing multi-view stereo pipeline, with straightforward generalization capability for different multi-view capture systems such as camera relative positioning and lenses. Given an initial depth estimation, CHOSEN iteratively re-samples and selects the best hypotheses, and automatically adapts to different metric or intrinsic scales determined by the capture system. The key to our approach is the application of contrastive learning in an appropriate solution space and a carefully designed hypothesis feature, based on which positive and negative hypotheses can be effectively distinguished. Integrated in a simple baseline multi-view stereo pipeline, CHOSEN delivers impressive quality in terms of depth and normal accuracy compared to many current deep learning based multi-view stereo pipelines.
Machine Learning, ICML
@@ 1 Introduction
Geometry acquisition through multi-view imagery is a crucial task in 3D computer vision. In the multi-view stereo matching (MVS) framework, image patches or features are matched and triangulated to find 3D points (Collins, 1996; Furukawa & Ponce, 2010). The dense geometry is usually represented as a depth map for each view. Due to view dependent appearance, occlusion and self-similarity in the real world scenarios, the matching signal is often too noisy to directly give accurate and complete geometry. Traditionally, this problem have been attacked by some carefully designed optimization and filtering scheme (Barron & Poole, 2016; Taniai et al., ), directly on the depth map or indirectly inside a matching cost volume, where they impose certain priors on the smoothness of the resulting geometry, making use of the confidence of matching and appearance in a larger spatial context.
With the advancement of deep learning, many research works have combined the ideas with traditional MVS. Early works (Yao et al., 2018, 2019) typically involve extracting convolutional features from the images, building one cost volume (indexed by a sequence monotone depth hypotheses $d_{i}$ for each pixel), or several in a multi-resolution hierarchy, and transform these cost volumes using 3D convolutions into a ”probability volume” so that depth map can be obtained by taking the expectation
@@
More recent works have more complicated designs regarding cost computation and aggregation (Wang et al., 2021), visibility and uncertainty estimation (Zhang et al., 2023a; Cheng et al., 2020), loss functions (Peng et al., 2022), feature backbones (Cao et al., 2023), feature fusion (Zhang et al., 2023b), and consequently have yielded impressive progress on various benchmarks. However, the evaluation has been limited to point cloud reconstruction, and little attention was paid to, and thus the lack thereof, the raw output’s quality in terms of depth and normal accuracy. Moreover, the fixed training and testing benchmarks also result in limited discussion on the learning and inference task under the variance of the multi-view acquisition setups, including different metric scales, camera relative positioning and lenses (focal lengths). Therefore, for a new capture setup substantially different from the training set, one is usually left with the job of determining what hyper-parameters to choose for cost volumes and probably retrain the model of choice. In fact, both of the problems can be attributed to the approach of using Eq.(1) to obtain depth estimations: depths hypotheses at different pixels, or across different datasets, may not be well aligned. The first results in non-smoothness, and the second results in poor generalization.
In this work, we turn our attention to the problem of learning to refine depth with multi-view images while keeping in mind that they can be from very different setups, and how to best sample hypotheses for depth refinement. We will discuss in Sec.3.1 the proper solution space to perform the learning task, and present an alternative to the approach of Eq.(1) in Sec.3.2. Much as inspired by the PatchMatch (Barnes et al., 2009; Bleyer et al., 2011) methodology, our framework, named CHOSEN, learns to refine the depth by iteratively re-sampling and selecting the best hypotheses, thus eliminating the need of a ”probability volume”.
@@
To demonstrate the effectiveness of CHOSEN, in Sec.3.3 we describe how we embed it in a minimalist MVS pipeline where the only extra learnable part is the feature extraction using light-weight U-Nets (Ronneberger et al., 2015). We perform comprehensive ablation experiments to justify our design, and compare the quality of our refined depths with various recent deep learning based MVS pipelines. Without bells and whistles, our baseline model is easy to train, fast to converge, and already delivers impressive depth estimation quality.
@@ 2 Related work
MVS algorithms are often categorized by the representation used to reconstruct the output scene, e.g. volume (Kutulakos & Seitz, 2000; Kar et al., 2017; Ji et al., 2017), point cloud (Furukawa & Ponce, 2010; Lhuillier & Quan, 2005; Lin et al., 2018; Insafutdinov & Dosovitskiy, 2018), depth map (Galliani et al., 2016; Tola et al., 2012; Schönberger et al., 2016), etc. However, depth map is probably the most flexible and efficient representation among existing ones. While depth map can be considered as a particular case of point cloud representation, (e.g., pixel-wise point), it reduces the reconstruction task to a per-view (2D) depth map estimation. These MVS approaches can be further grouped in hand-crafted (traditional) methods or learning based solutions. Traditional MVS pipelines extend the two-view case (Scharstein & Szeliski, 2002) by introducing a view-selection step that aggregates the cost from multiple images to a given reference view. The view selection can be performed per-camera (Galliani et al., 2015) or per-pixel (Schönberger et al., 2016). These approaches (Galliani et al., 2015, 2016; Tola et al., 2012; Schönberger et al., 2016) then rely on well engineered photometric cost functions (ZNCC, SSD, SAD, etc.) to estimate the scene 3D geometry, by selecting the best depth hypothesis that leads to the lowest aggregated cost. However these cost functions usually perform poorly on texture-less or occluded areas, and under complex lighting environments where photometric consistency is unreliable. Hence further post-processing and propagation steps are used to improve the final estimate (Schönberger et al., 2016; Galliani et al., 2015). We refer readers to (Furukawa & Hernández, 2015) for additional details regarding traditional multi-view stereo pipelines.
@@
Recently, MVS algorithms have showed impressive quality of 3D reconstructions in terms of accuracy and completeness, mostly thanks to the increase popularity of deep learning based solutions (Yao et al., 2018, 2019; Yu & Gao, 2020; Gu et al., 2020; Wang et al., 2021; Zhang et al., 2023b; Cao et al., 2023). These methods often make use of multi-scale feature extractors (Cheng et al., 2020; Tankovich et al., 2021; Chang & Chen, 2018), cost-volumes (Kendall et al., 2017; Khamis et al., 2018; Zhang et al., 2018), and guided refinement (Yao et al., 2018; Pang et al., 2017; Khamis et al., 2018) to retrieve the final 3D estimate. Typically, they leverage U-Nets (Ronneberger et al., 2015) to build a single or a hierarchy of cost volumes with predetermined depth hypotheses. Then, these cost volumes are regularized using 3D convolutions and the final depth map is regressed from the regularized probability volume. However, to achieve high resolution depth accuracy, it requires sampling a large number of depth planes, which is limited by memory consumption.
@@
Researchers are also opening other frontiers in deep learning based MVS. UCS-Net (Cheng et al., 2020) and Vis-MVSNet (Zhang et al., 2023a) use some uncertainty estimate for an adaptive generation of cost volumes. PVSNet (Xu & Tao, 2020) and PatchmatchNet (Wang et al., 2021) learn to predict visibility for each source image. The approach of Eq.(1) is considered critically in (Peng et al., 2022), although it was only considered from the loss function perspective. GeoMVSNet (Zhang et al., 2023b) proposes a geometry fusion mechanism in the MVS pipline. Most recently, transformer based methods (Ding et al., 2022; Cao et al., 2023, 2024) exploits the attention mechanism for more robust matching and context awareness. In particular, MVSFormer (Cao et al., 2023) has combined this approach with the powerful pre-trained DINO features (Oquab et al., 2023).
@@
It is worth mentioning that existing methods usually need to hand-tune hyper-parameters such as the depth range, hypotheses spacing, and number of hypotheses to ensure sufficient and accurate coverage for the new application. Some of these methods adopt 2D convolutional neural networks to obtain final depth estimations, using RGB image to guide depth up-sampling and refinement (Khamis et al., 2018; Hui et al., 2016; Wang et al., 2021). Consequently, these methods often generalize poorly to new camera setups or new scenes.
@@ 3 Framework
Figure 1: CHOSEN’s hypotheses sampling and ranking mechanism. Each hypothesis of each pixel is processed independently.
@@ 3.1 Hypothesis representation and sampling
Inspired by the PatchMatch framework (Barnes et al., 2009; Bleyer et al., 2011), we use spatial hypotheses to expand good solutions into their vicinity. Specifically, we sample a set of hypotheses for each pixel by collecting the propagated depths from its spatial neighbors. The propagation is facilitated through first order approximation. Denote $\partial d=(\partial_{x}d,\partial_{y}d)$ to be gradient of the depth $d$, $\mathbf{p}^{\prime}=\mathbf{p}+\Delta\mathbf{p}$ be a spatial neighbor of the pixel $\mathbf{p}$. The propagated depth from $\mathbf{p}^{\prime}$ to $\mathbf{p}$ is defined as
@@ 3.2 Contrastive learning for hypothesis ranking
Finally, $\mathbf{f}(\hat{d})$ is a learned context feature of the previous refined depth, whose inputs are the concatenation of $\mathbf{c}(\hat{d})$, $\mathbf{e}(\hat{d})$ and the feature from a U-Net describing the reference view’s high level appearance. We use a two-layer convolution network to initialize $\mathbf{f}(\hat{d})$, and subsequently we use a convolutional gated recurrent unit (GRU) to update $\mathbf{f}(\hat{d})$ given the updated $\hat{d}$. We remark that the overall iterative update is similar to the RAFT (Teed & Deng, 2020) framework, with the addition of the geometric error term $\mathbf{e}(d)$. We will discuss its crucial contribution in our ablation study in Sec.4.2.
@@ 3.3 Baseline MVS design
It should be noted that our method can be embedded end-to-end in any deep learning based MVS framework. Here we will demonstrate the design of a simple end-to-end MVS pipeline. We use simple feature correlation as the matching cost, so that there are no learnable components for cost volume filtering, depth refinement, or confidence prediction. The only learnable parts are the ranking MLP, context update convolutional layers, the light-weight U-Nets used for extracting matching feature as well as the high level appearance of the reference view.
Hence our matching component is very noisy, and suboptimal in the cases where there are large view distortion, relative rotation or scale change. One should expect straightforward improvement when additional techniques are incorporated, such as cost volume filtering (Yao et al., 2018), learnable aggregation (Zhang et al., 2023a), other feature extraction backbones (Cao et al., 2023), etc. Even with noisy feature correlation matching signal, we show in Sec.4.3 that it can achieve superior quality in depth and normal compared to many state-of-art MVS works on the DTU dataset, highlighting the robustness of our approach.
@@
We use a two-stream feature extraction architecture similar to (Teed & Deng, 2020). First, we use a matching feature network extracts distinctive features that are then used to evaluate the cost of depth hypotheses between the reference and other source views. Second, we use a context feature network that learns high-level spatial cues for learnable hypothesis selection, applied only on the reference image. Both networks are based on a light-weight U-Net architecture (Ronneberger et al., 2015) which extracts features from coarse to fine.
@@
Similar to the contrastive loss used in (Tankovich et al., 2021), we supervise the feature matching using a simple contrastive loss
@@ 4 Experiments
MVSFormer (Cao et al., 2023) $61.92$ $0.4248$ $18.96$ $44.17$
GeoMVSNet (Zhang et al., 2023b) $63.93$ $0.4089$ $15.18$ $38.40$
GBi-Net (Mi et al., 2022) $34.01$ $0.4531$ $13.78$ $33.41$
IterMVS (Wang et al., 2022) $51.09$ $0.4524$ $16.67$ $37.50$
PatchMatchNet (Wang et al., 2021) $56.19$ $0.4260$ $18.03$ $41.82$
UCSNet (Gu et al., 2020) $59.03$ $0.3942$ $31.31$ $54.39$
Table 1: Comparison for estimated depths and normals on DTU testing set. We mark our best results separately for $\nicefrac{{1}}{{8}}$ and $\nicefrac{{1}}{{4}}$ output resolutions. Notice that so long as the finest resolution matching features are used, the final depth accuracy metrics for these two output resolutions are very similar. Our simple baselines offer significant improvement in terms of depth quality even compared to the strongest state-of-art MVS pipelines.
@@ 4.1 Implementation details
The model is trained and tested on a single NVIDIA A100 GPU (40G). We first train with batch size 4, at input resolution $600\times 800$, up to 200k iterations using the default Adam optimizer (Kingma & Ba, 2014) at a learning rate of $0.001$. We then fine-tune for up to 50k iterations at input resolution $1200\times 1600$ at a learning rate of $0.0005$. During training, we fix the closest source view and randomly sample two from the remaining source views, which is the only data augmentation we used. For testing, it takes around $0.6$ seconds to run at $1200\times 1600$ with $5$ source views. We provide results both from models trained on DTU (Jensen et al., 2014) training set only, and on DTU mixed with the BlendedMVS (Yao et al., 2020) dataset, which will be tagged specifically.
@@
  • %$\%<x^{\circ}$(@$<y$mm): the percentage of pixels where the normal is within $x^{\circ}$ of angular error out of all pixels whose absolute error is less than $y$-mm.
@@
Figure 4: Direct application of our DTU + BlendedMVS trained baseline model on instances from MultiFace (Knapitsch et al., 2017), Tanks & Temples and ETH3D (Schops et al., 2017) datasets. Our simple baseline achieves consistent generalization ability even though trained on substantially different data.

### v2

@@ Abstract
We propose CHOSEN, a simple yet flexible, robust and effective multi-view depth refinement framework. It can provide significant improvement in depth and normal quality, and can be integrated in existing multi-view stereo pipelines with minimal modifications. Given an initial depth estimation, CHOSEN iteratively re-samples and selects the best hypotheses. The key to our approach is the application of contrastive learning in an appropriate solution space and a carefully designed hypothesis feature, based on which positive and negative hypotheses can be effectively distinguished. We integrated CHOSEN in a basic multi-view stereo pipeline, and show that it can deliver impressive quality in terms of depth and normal accuracy compared to many other top deep learning based multi-view stereo pipelines.
@@ 1 Introduction
Geometry acquisition through multi-view imagery is a crucial task in 3D computer vision. In the multi-view stereo matching (MVS) framework, image patches or features are matched and triangulated to find 3D points [9, 12]. The dense geometry is usually represented as a depth map for each view. Due to view dependent appearance, occlusion and self-similarity in the real world scenarios, the matching signal is often too noisy to directly give accurate and complete geometry. Traditionally, this problem have been attacked by some carefully designed optimization and filtering scheme [3, 36], directly on the depth map or indirectly inside a matching cost volume, where they impose certain priors on the smoothness of the resulting geometry, making use of the confidence of matching and appearance in a larger spatial context.
Figure 1: CHOSEN’s hypotheses sampling and ranking mechanism. Assuming an initial depth estimation, we gather depth hypotheses from its perturbations or its spatial neighbors. For each hypothesis, we combine the matching cost, a second order smoothness term and a context feature to form the input of a learned score function represented by an MLP. The one with the highest score will be selected to update refined depth. The process is performed iteratively.
With the advancement of deep learning, many research works have combined the ideas with traditional MVS. Early works [44, 45] typically involve extracting convolutional features from the images, building one cost volume (indexed by a sequence monotone depth hypotheses $d_{i}$ for each pixel), or several in a multi-resolution hierarchy, and transform these cost volumes using 3D convolutions into a ”probability volume” so that depth map can be obtained by taking the expectation
@@
More recent works have more complicated designs regarding cost computation and aggregation [40], visibility and uncertainty estimation [49, 8], loss functions [31], feature backbones [5], feature fusion [51], and consequently have yielded impressive progress on various benchmarks. The evaluation has been limited to point cloud reconstruction, and little attention was paid to, and thus the lack thereof, the raw output’s quality in terms of depth and normal accuracy. Looking closely at the predicted depth and its associated normals, we can often find quantization-like artifacts. Moreover, the fixed training and testing benchmarks also result in limited discussion on the learning and inference task under the variance of the multi-view acquisition setups, including different metric scales, camera relative positioning and lenses (focal lengths). Therefore, for a new capture setup substantially different from the training set, one may need to determine what hyper-parameters to use, especially the number of depth proposals to sample from a certain depth range, and retrain the model of choice to get the optimal performance.
To some extent, the problem can be attributed to the approach of using Eq.(1) to obtain depth estimations: the distribution of depths hypotheses at different pixels, or across different datasets, can be arbitrary. It is thus difficult to guarantee the learned probability weights to produce consistent results: in a smooth spatial neighborhood, small perturbations in the ${d_{i}}$ results in non-smoothness in the predicted depth and normal; across different datasets, the inappropriately spaced ${d_{i}}$ that are too sparse or too coarse results in poor generalization.
In this work, we turn our attention to the problem of learning to refine depth with multi-view images while keeping in mind that they can be from very different setups, and how to best sample hypotheses for depth refinement. We will discuss in Sec.3.1 the proper solution space to perform the learning task, and present an alternative to the approach of Eq.(1) in Sec.3.2. Much as inspired by the PatchMatch [2, 4] methodology, our framework, named CHOSEN, learns to refine the depth by iteratively re-sampling and selecting the best hypotheses, thus eliminating the need of a ”probability volume”.
@@
To demonstrate the effectiveness of CHOSEN, in Sec.3.3 we describe how we embed it in a minimalist MVS pipeline where the only extra learnable part is the feature extraction using light-weight U-Nets [32]. We perform comprehensive ablation experiments to justify our design, and compare the quality of our refined depths with various recent deep learning based MVS pipelines. Without bells and whistles, our baseline model is easy to train, fast to converge, and already delivers impressive depth estimation quality.
@@ 2 Related work
MVS algorithms are often categorized by the representation used to reconstruct the output scene, e.g. volume [25, 20, 19], point cloud [12, 26, 27, 17], depth map [14, 39, 34], etc. However, depth map is probably the most flexible and efficient representation among existing ones. While depth map can be considered as a particular case of point cloud representation, (e.g., pixel-wise point), it reduces the reconstruction task to a per-view (2D) depth map estimation. These MVS approaches can be further grouped in hand-crafted (traditional) methods or learning based solutions. Traditional MVS pipelines extend the two-view case [33] by introducing a view-selection step that aggregates the cost from multiple images to a given reference view. The view selection can be performed per-camera [13] or per-pixel [34]. These approaches [13, 14, 39, 34] then rely on well engineered photometric cost functions (ZNCC, SSD, SAD, etc.) to estimate the scene 3D geometry, by selecting the best depth hypothesis that leads to the lowest aggregated cost. However these cost functions usually perform poorly on texture-less or occluded areas, and under complex lighting environments where photometric consistency is unreliable. Hence further post-processing and propagation steps are used to improve the final estimate [34, 13]. We refer readers to [11] for additional details regarding traditional multi-view stereo pipelines.
@@
Recently, MVS algorithms have showed impressive quality of 3D reconstructions in terms of accuracy and completeness, mostly thanks to the increase popularity of deep learning based solutions [44, 45, 47, 15, 40, 51, 5]. These methods often make use of multi-scale feature extractors [8, 37, 7], cost-volumes [21, 22, 50], and guided refinement [44, 30, 22] to retrieve the final 3D estimate. Typically, they leverage U-Nets [32] to build a single or a hierarchy of cost volumes with predetermined depth hypotheses. Then, these cost volumes are regularized using 3D convolutions and the final depth map is regressed from the regularized probability volume. However, to achieve high resolution depth accuracy, it requires sampling a large number of depth planes, which is limited by memory consumption.
@@
Researchers are also opening other frontiers in deep learning based MVS. UCS-Net [8] and Vis-MVSNet [49] use some uncertainty estimate for an adaptive generation of cost volumes. PVSNet [43] and PatchmatchNet [40] learn to predict visibility for each source image. The approach of Eq.(1) is considered critically in [31], although it was only considered from the loss function perspective. GeoMVSNet [51] proposes a geometry fusion mechanism in the MVS pipline. Most recently, transformer based methods [10, 5, 6] exploits the attention mechanism for more robust matching and context awareness. In particular, MVSFormer [5] has combined this approach with the powerful pre-trained DINO features [29].
@@
It is worth mentioning that existing methods usually need to hand-tune hyper-parameters such as the depth range, hypotheses spacing, and number of hypotheses to ensure sufficient and accurate coverage for the new application. Some of these methods adopt 2D convolutional neural networks to obtain final depth estimations, using RGB image to guide depth up-sampling and refinement [22, 16, 40]. Consequently, these methods often generalize poorly to new camera setups or new scenes.
@@ 3.1 Hypothesis representation and sampling
Inspired by the PatchMatch framework [2, 4], we use spatial hypotheses to expand good solutions into their vicinity. Specifically, we sample a set of hypotheses for each pixel by collecting the propagated depths from its spatial neighbors. The propagation is facilitated through first order approximation. Denote $\partial d=(\partial_{x}d,\partial_{y}d)$ to be gradient of the depth $d$, $\mathbf{p}^{\prime}=\mathbf{p}+\Delta\mathbf{p}$ be a spatial neighbor of the pixel $\mathbf{p}$. The propagated depth from $\mathbf{p}^{\prime}$ to $\mathbf{p}$ is defined as
@@ 3.2 Contrastive learning for hypothesis ranking
Finally, $\mathbf{f}(\hat{d})$ is a learned context feature of the previous refined depth, whose inputs are the concatenation of $\mathbf{c}(\hat{d})$, $\mathbf{e}(\hat{d})$ and the feature from a U-Net describing the reference view’s high level appearance. We use a two-layer convolution network to initialize $\mathbf{f}(\hat{d})$, and subsequently we use a convolutional gated recurrent unit (GRU) to update $\mathbf{f}(\hat{d})$ given the updated $\hat{d}$. We remark that the overall iterative update is similar to the RAFT [38] framework, with the addition of the geometric error term $\mathbf{e}(d)$. We will discuss its crucial contribution in our ablation study in Sec.4.2.
@@
As result, the ranking module will learn to put most of the weight on the set $\mathcal{D}^{+}$. We found this formulation most effective in training, and much superior compared to learning to put all weight on the closest sample to the ground truth. We believe it’s due to that during spatial propagation, many samples will be already close to the ground truth and supervising on the closest sample introduces unnecessary competition among good samples.
Figure 3: Evolution of the refined depth in our baseline MVS model. The first row shows the ground truth and input images, and we mark each row with the cost volume pyramid configuration. As one can see, the winner-take-all initialization from the cost volume is usually very noisy, but nevertheless contain some accurate estimations. By iteratively re-sample and selecting the best hypotheses, the depth and quality are significantly improved.
@@ 3.3 Baseline MVS design
Here we will demonstrate the design of a simple end-to-end MVS pipeline with our depth refinement method. We have deliberately designed it such that there are no learnable components for cost volume construction, depth refinement, or confidence prediction. The only learnable parts are the ranking MLP, context update convolutional layers, the light-weight U-Nets used for extracting matching feature as well as the high level appearance of the reference view. Hence our cost volume construction should be inferior and more noisy compared to many existing MVS models that are equipped with cost volume filtering. Even with our simple pipeline, we show in Sec.4.3 that it can achieve much more superior quality in depth and normal compared to many state-of-art MVS works on the DTU dataset.
@@
We use a two-stream feature extraction architecture similar to [38]. First, we use a matching feature network extracts distinctive features that are then used to evaluate the cost of depth hypotheses between the reference and other source views. Second, we use a context feature network that learns high-level spatial cues for learnable hypothesis selection, applied only on the reference image. Both networks are based on a light-weight U-Net architecture [32] which extracts features from coarse to fine.
@@
Similar to the contrastive loss used in [37], we supervise the feature matching using a simple contrastive loss
@@
3.4 Integration in MVS pipeline
It’s clear to see that CHOSEN can be embedded end-to-end in existing deep learning based MVS framework with little modification needed, since CHOSEN only needs the cost, second order error term and context feature to rank the depth hypotheses. One may substitute our simplistic yet noisy cost volume construction in the baseline with the techniques proposed in recent works, such as cost volume filtering [44] and other feature backbones [5], etc. The main thing to consider for the integration is that the depth hypotheses must be converted to the pseudo disparity representation in order to be processed by CHOSEN. Since the spacing of the hypotheses in the cost volume is crucial for contrastive learning, it might be necessary to retrain the model with CHOSEN from scratch since the original model might be sensitive to different spacing. In Sec4.5 We illustrate the some results demonstrating the further improvement CHOSEN by introducing cost volume filtering based on MVSNet [44].
@@ 4 Experiments
MVSFormer [5] $61.92$ $0.4248$ $18.96$ $44.17$
GeoMVSNet [51] $63.93$ $0.4089$ $15.18$ $38.40$
GBi-Net [28] $34.01$ $0.4531$ $13.78$ $33.41$
IterMVS [41] $51.09$ $0.4524$ $16.67$ $37.50$
PatchMatchNet [40] $56.19$ $0.4260$ $18.03$ $41.82$
UCSNet [15] $59.03$ $0.3942$ $31.31$ $54.39$
Table 1: Comparison for estimated depths and normals on DTU testing set. The metric is computed on all the valid ground truth pixels. We mark our best results separately for $\nicefrac{{1}}{{8}}$ and $\nicefrac{{1}}{{4}}$ resolutions. The results from other methods are evaluated at $\nicefrac{{1}}{{4}}$ resolutions with nearest neighbor down-sampling. Notice that so long as the finest resolution matching features are used, the final depth accuracy metrics for these two output resolutions are very similar. Our simple baselines offer significant improvement in terms of depth quality even compared to the strongest state-of-art MVS pipelines.
@@ 4.1 Implementation details
Our baseline model is trained and tested on a single NVIDIA A100 GPU (40G). The total trainable parameter count is about 1.1 million, including the matching and context U-Nets. Our selection module’s learnable parameter count is about 781k. We first train with batch size 4, at input resolution $600\times 800$, up to 200k iterations using the default Adam optimizer [23] at a learning rate of $0.001$. We then fine-tune for up to 50k iterations at input resolution $1200\times 1600$ at a learning rate of $0.0005$. During training, we fix the closest source view and randomly sample two from the remaining source views, which is the only data augmentation we used. For testing, it takes around $0.6$ seconds to run at $1200\times 1600$ with $5$ source views, with a total of 12 iterations of hypotheses ranking (4 iterations for each cost volume config). We provide results both from models trained on DTU [18] training set only, and on DTU mixed with the BlendedMVS [46] dataset, which will be tagged specifically.
@@
  • %$\%<x^{\circ}$(@$<y$-dsp): the percentage of pixels where the normal is within $x^{\circ}$ of angular error out of all pixels whose absolute error is less than $y$-pseudo disparity.
@@
Figure 5: Direct application of our DTU + BlendedMVS trained baseline model on instances from MultiFace [24], Tanks & Temples and ETH3D [35] datasets. Our simple baseline achieves consistent generalization ability even though trained on substantially different data.
Figure 6: From left to right: The reference view, CHOSEN in baseline MVS, and CHOSEN in baseline w/ 3D-CNN. 3D-CNN is applied to the full range cost volume. The improvement is most significant in large texture-less or repetitive regions.